Find a matnx A such that X3 (#1) Let X= 3x2 x 6=XAx (x)= In addition, venfy thet ag x 2Ax holds.
Answer #1

X- 2 2. XAX 0x)=3X4tX3-7X-63 3 XI XI xAX 4- -3 2. X 3 -3 X3 XI 3 ХАX -О e 3 X2 4 2- Хз -3 Savice X2 X3 (ca = xA X From we canSarice xA A X Әх Swne (xA)= Ax A X AX fe ax Sica A A - AX AX (K)be ai = 2A X e Әх Aus uee